What do you mean by recoil of gun? Obtain an expression for recoil velocity? A shell of mass 30g is fired from a gun, with a speed of 80m/s. Calculate the recoil velocity of gun if its mass is 200g.


1
Expert's answer
2020-12-25T14:07:03-0500

"\text{A recoil is a movement backwards, usually from some force or impact. }"

"\text{We use the law of conservation of momentum:}"

"m_g*V_g+m_b*V_b= const"

"\text{where:}"

"m_g,V_g-\text{gun mass and speed}"

"m_b,V_b-\text{bullet mass and speed}"

"const = 0 \text{ since before the start of the movement the speed of the gun }"

"\text{and the bullet are 0}"

"m_g*V_g+m_b*V_b= 0"

"V_g= -\frac{m_b*V_b}{m_g}"

"V_g\text{ has a minus sign because it is opposite to the direction of the bullet}"

"|V_g|= \frac{0.03*80}{0.2}=12m/s"

Answer:12m/s gun speed